im really lost with this. A(-2;-2), B, C(4;4) and D are the vertices of a rhombus. Find the the equation of diagnol BD. the workout i cant complete im totally lost . i would appreciate it if u could heelp me with this plz
Answer by scott8148(6628)

the diagonals of a rhombus are perpendicular and they bisect each other

this means that the slope of BD is the negative reciprocal of the slope of AC
__ and BD passes through the midpoint of AC

slope of AC __ m=(4-(-2))/(4-(-2)) __ m=1
__ so the slope of BD is -1

midpoint of AC __ ((-2+4)/2,(-2+4)/2) __ (1,1)

using point-slope for BD __ y-1=-1(x-1)
__ solving for y (slope-intercept) __ y=-x+2